OpenStudy (anonymous):

determine where the function f(x) is continuous (x^2+7)/(x^2+x-6)

OpenStudy (nory):

The question here is where is it _not_ continuous?

OpenStudy (nory):

That's at the points where it's undefined, or makes a jump.

OpenStudy (nory):

This function has undefined points. Can you see why? It's because the denominator is going to be 0 sometimes.

OpenStudy (anonymous):

If that is a rational function (a fraction) then we find where the denominator is equal to zero. Whenever the denominator is equal to 0 we have a discontinuity. That is the function is not defined because we cannot divide by the number 0.

OpenStudy (anonymous):

so this is not ever going to have the denominator = zero?

OpenStudy (anonymous):

So we solve for x^2+x-6=0 we can use the quadratic equation or factor it.

OpenStudy (anonymous):

Exactly. Thus, we are looking for when the the denominator is equal to zero and exclude (do not include) this for the domain (allowed values of x).

OpenStudy (anonymous):

Can you solve for x^2+x-6=0?

OpenStudy (anonymous):

so basically this is never continuous?

OpenStudy (anonymous):

Not quite. This function is continuos everywhere except when x^2+x-6=0 (i.e whichever x values satisfy the equation x^2+x-6=0)

OpenStudy (anonymous):

-3,2

OpenStudy (anonymous):

exactly!!!

OpenStudy (anonymous):

:-) awesome

OpenStudy (anonymous):

So we conclude: the function f(x) is continuous everywhere except when x=-3 and x=2

OpenStudy (anonymous):

hmm, this concept confuses me! what happens to the numerator? it doesn't matter?

OpenStudy (anonymous):

this is a better way to look at it. When does a fraction make sense? a/b makes sense whenever b is not 0. a can be any number, since we can divide any number by another (except for 0)

OpenStudy (anonymous):

where the function is continuous means for which x-values does our function defined (for which x-values can we evaluate our function)...we cannot evaluate the function when we divide by 0 i.e. when the denominator is 0 b/c we don't know how to divide by 0

OpenStudy (anonymous):

***is our function defined (not "does our function defined)

OpenStudy (anonymous):

exactly the numerator can equal anything because whichever values of x we plug in we can evaluate it

OpenStudy (anonymous):

hmm okay! I hope I can burn this concept into my brain!